OpenStudy (anonymous):

What is the arc length of the arc subtended in a circle with radius 6 and an angle of 7 8 ?

OpenStudy (amistre64):

consider arclength so be a division of circumference

OpenStudy (anonymous):

ok thanks

OpenStudy (amistre64):

since the circumference of a circle is simply: 2pi r the percentage of that which is related to \(\theta\) degrees is: \(\theta\)7 out of 360, or simply a multipler of \(\theta/360\) \[\frac{\theta}{360}~2\pi~r\]

OpenStudy (anonymous):

oh ok i see now thanks